# Entropy Functions on Two-Dimensional Faces of Polymatroidal Region of Degree Four: Part II: Information Theoretic Constraints Breed New Combinatorial Structures

### v1

@@ III-C Entropy functions on faces involving Latin square decompositions
$H(X_{1})=H(\dfrac{\alpha_{0}}{v^{2}},\dfrac{\alpha_{1}}{v^{2}},\dots,\dfrac{\alpha_{t-1}}{v^{2}}),$ (III.3.70)
where $\alpha_{i}$ denotes the times of the $i\in\mathbb{I}_{t}$ that occurs in $\mathbf{T}(1)$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=2a+b.$ (III.3.71)
$H(X_{2})$ $=H(X_{3})=H(X_{4})=a+b$ (III.3.72)
Note that $a+b=\log v$, we obtain tat
$a=H(\dfrac{\alpha_{0}}{v^{2}},\dfrac{\alpha_{1}}{v^{2}},\dots,\dfrac{\alpha_{t-1}}{v^{2}})-\log v.$ (III.3.73)
@@ III-D Entropy functions on the face $(\hat{U}_{3,5}^{4},U_{2,4})$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i<j,i,j\in N_{4},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),|K|=2,\{4\}\subseteq K\subseteq N_{4}.$
@@
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{4})p(x_{1},x_{2},x_{3},x_{4}),$ (III.4.11)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1},x_{2},x_{3},x_{4}),$ (III.4.12)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{3},x_{4})p(x_{1},x_{2},x_{3},x_{4}).$ (III.4.13)
By (III.4.1), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (III.4.11), we have
@@
It can be seen that each of such constructed $\mathbf{T}_{i}$ is a $\mathrm{VOA}(U_{2,4},v)$. Let $(X_{i},i\in N_{4})$ be distributed on the rows of $\mathbf{T}\triangleq\left[\begin{matrix}\mathbf{T}_{0},\\ \dots\\ \mathbf{T}_{v-1}\end{matrix}\right]$ such that the probability mass of each row of $\mathbf{T}_{i}$ is $\frac{p_{i}}{v^{2}}$. Now we show the entropy function of $(X_{i},i\in N_{4})$ is $(a,b)$. By (III.4.24), $|\mathcal{X}_{4}|=v^{2}$ and each entry $j\in\mathcal{X}_{4}$ occurs only in one of the arrays $\mathbf{T}_{0}(4)$, $\dots$, $\mathbf{T}_{v-1}(4)$, which implies that for any subset $A$ such that $\{4\}\subseteq A\subseteq N_{4}$, $x_{A}\in\mathcal{X}_{A}$ will appear
$v^{\mathbf{r}_{U_{2,4}}(N_{4})-\mathbf{r}_{U_{2,4}}(A)}$ (III.4.25)
times where $\mathbf{r}_{U_{2,4}}$ is the rank function of $U_{2,4}$. So
$H(X_{A})$ $=H(\underbrace{\frac{p_{0}v^{\mathbf{r}_{U_{2,4}}(N_{4})-\mathbf{r}_{U_{2,4}}(A)}}{v^{2}},\dots,\frac{p_{0}v^{\mathbf{r}_{U_{2,4}}(N_{4})-\mathbf{r}_{U_{2,4}}(A)}}{v^{2}}}_{v^{\mathbf{r}_{U_{2,4}}(A)}},\dots,\underbrace{\frac{p_{v-1}v^{\mathbf{r}_{U_{2,4}}(N_{4})-\mathbf{r}_{U_{2,4}}(A)}}{v^{2}},\dots,\frac{p_{v-1}v^{\mathbf{r}_{U_{2,4}}(N_{4})-\mathbf{r}_{U_{2,4}}(A)}}{v^{2}}}_{v^{\mathbf{r}_{U_{2,4}}(A)}})$ (III.4.26)
$=H(p_{0},p_{1},\dots,p_{v-1})+\log{v}\cdot\mathbf{r}_{U_{2,4}}(A)=a+(a+b)\cdot\mathbf{r}_{U_{2,4}}(A)=a(1+\mathbf{r}_{U_{2,4}}(A))+b\mathbf{r}_{U_{2,4}}(A).$ (III.4.27)
@@
$H(X_{3})=H(\frac{v(p_{0}+p_{1}+\dots,+p_{v-1})}{v^{2}},\dots,\frac{v(p_{0}.+p_{1}+\dots,+p_{v-1})}{v^{2}})=\log{v}.$ (III.4.28)
@@
$H(X_{1},X_{3})=H(X_{2},X_{3})=H(\frac{p_{0}+p_{1}+\dots,+p_{v-1}}{v^{2}},\dots,\frac{p_{0}+p_{1}+\dots,+p_{v-1}}{v^{2}})=2\log{v}.$ (III.4.29)
@@
$H(X_{1},X_{2},X_{3})$ $=H(\underbrace{\frac{p_{0}}{k^{2}},\frac{p_{0}}{v^{2}},\dots,\frac{p_{0}}{v^{2}}}_{v^{2}},\dots,\underbrace{\frac{p_{v-1}}{v^{2}},\frac{p_{v-1}}{v^{2}},\dots,\frac{p_{v-1}}{v^{2}}}_{v^{2}})$ (III.4.30)
$=H(p_{0},\dots,p_{v-1})+2\log{v}=a+2(a+b)=3a+2b.$ (III.4.31)
@@
$p_{1}=p_{4}=p_{6}=p_{7},$ (III.4.32)
$p_{2}=p_{3}=p_{5}=p_{8}.$ (III.4.33)
Assume that either (III.4.32) or (III.4.33) vanishes, then $\mathbf{T}_{0}$ degenerates a $\mathrm{VOA}(U_{2,3},2)$ $\mathbf{T}_{1}$ and $(X_{1},X_{2},X_{3})$ is uniformly distributed on the rows of $\mathbf{T}_{1}$. Together with (III.4.18) and (III.4.21),
$a=0,b=\log 2,$ (III.4.34)
which contradicts the fact that $\mathbf{h}=(0,\log 2)$ is non-entropic. Hence, both (III.4.32) and (III.4.33) must be positive. By (III.4.1), $X_{4}$ is a function of $(X_{1},X_{2},X_{3})$, which implies that $(X_{i},i\in N_{4})$ must be distributed on a $\mathbf{T}$ such that $\mathbf{T}(1,2,3)=\mathbf{T}_{0}$. By (III.4.15) and (III.4.16), for each $x_{4}\in\mathcal{X}_{4},x_{j}\in\mathcal{X}_{j}$, $(x_{j},x_{4})$ appear exactly once on the row of $\mathbf{T}(j,4)$ for $j=1,2,3$. Additionally, $X_{4}$ is independent of $X_{i},i=1,2,3$ by (III.4.7), (III.4.9), and (III.4.10). There exists a unique $\mathbf{T}$ satisfying the above information equalities up to symmetry, and
@@
Now we show an inner bound on the entropy region on $F$ that $a+b=\log 6$ and $\log 2\leq a\leq\log 6$. Let $p_{i}>0$, $i=0,1,2$, and $p_{0}+p_{1}+p_{2}=1$ and $H(p_{0},p_{1},p_{2})=a-\log 2$. Let $\mathbf{T}^{(k)}_{0},k=1,2$ be arrays as follows.
@@
$\mathbf{T}^{(k)}_{i}(1,2)$ $=\mathbf{T}^{(k)}_{1}(1,2),$ (III.4.40)
$\mathbf{T}^{(k)}_{i}(3;j)$ $=\sigma_{i}(\mathbf{T}^{(k)}_{1}(3;j))\text{ for }j\in N_{36}$ (III.4.41)
$\mathbf{T}^{(k)}_{i}(4;j)$ $=\mathbf{T}^{(k)}_{1}(4;j)+12i\text{ for }j\in N_{36},$ (III.4.42)
@@
$H(X_{4})$ $=H(\underbrace{\frac{p_{0}}{12},\frac{p_{0}}{12},\dots,\frac{p_{0}}{12}}_{12},\underbrace{\frac{p_{1}}{12},\frac{p_{1}}{12},\dots,\frac{p_{1}}{12}}_{12},\underbrace{\frac{p_{2}}{12},\frac{p_{2}}{12},\dots,\frac{p_{2}}{12}}_{12})$ (III.4.44)
$=H(p_{0},p_{1},p_{2})+\log{12}=a-\log 2+\log 12=a+\log{6}=2a+b.$ (III.4.45)
@@
$p_{X_{1}}(j)=\frac{12p_{0}}{72}+\frac{12p_{1}}{72}+\frac{12p_{2}}{72}=\frac{1}{6},$ (III.4.46)
@@
$H(X_{1},X_{4})$ $=H(X_{1},X_{2},X_{3},X_{4})$ (III.4.49)
$=H(\underbrace{\frac{p_{0}}{72},\frac{p_{0}}{72},\dots,\frac{p_{0}}{72}}_{72},\underbrace{\frac{p_{1}}{72},\frac{p_{1}}{72},\dots,\frac{p_{1}}{72}}_{12},\underbrace{\frac{p_{2}}{72},\frac{p_{2}}{72},\dots,\frac{p_{2}}{72}}_{72})$ (III.4.50)
$=H(p_{0},p_{1},p_{2})+\log{72}=a-\log 2+\log 72=a+\log{36}=3a+2b.$ (III.4.51)
@@
$p_{X_{1}X_{3}}(x_{1},x_{2})=\frac{2p_{0}}{72}+\frac{2p_{1}}{72}+\frac{2p_{2}}{72}=\frac{1}{36},$ (III.4.54)
@@ III-E Discussion
In this section, we characterized $10$ types of $2$-dimensional faces of $\Gamma_{4}$. As each $\mathrm{MVOA}(P,v)$ with $M$ a rank 2 or greater integer polymatroid corresponds to a type of orthogonal Latin hypercubes, the characterization of a face with both extreme rays containing an integer polymatroid exceeding 1, in Subsection III-B-III-D breeds a new combinatorial structure which can be considered as an intermediate form of the two types of orthogonal Latin hypercubes. Specifically, in Subsection III-C, three faces are characterized by the Latin square decompositions, which can be considered as three new types of Latin square substructures.

### v2

@@ III-C Entropy functions on faces involving Latin square decompositions
$H(X_{1})=H(\dfrac{\alpha_{0}}{v^{2}},\dfrac{\alpha_{1}}{v^{2}},\dots,\dfrac{\alpha_{t-1}}{v^{2}}),$ (III.3.71)
where $\alpha_{i}$ denotes the times of $i\in\mathbb{I}_{t}$ that occurs in $\mathbf{T}(1)$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=2a+b.$ (III.3.72)
$H(X_{2})$ $=H(X_{3})=H(X_{4})=a+b$ (III.3.73)
Note that $a+b=\log v$, we obtain that
$a=H(\dfrac{\alpha_{0}}{v^{2}},\dfrac{\alpha_{1}}{v^{2}},\dots,\dfrac{\alpha_{t-1}}{v^{2}})-\log v.$ (III.3.74)
@@ III-D Entropy functions on the face $(\hat{U}_{3,5}^{4},U_{2,4})$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i,j\in N_{4},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),$
$\qquad\quad|K|=2,\{4\}\subseteq K\subseteq N_{4}.$
@@
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}),$ (III.4.11)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1234}),$ (III.4.12)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{3},x_{4})p(x_{1234}).$ (III.4.13)
By (III.4.1), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1234})$ on either side of (III.4.11), we have
@@
It can be seen that each of such constructed $\mathbf{T}_{i}$ is a $\mathrm{VOA}(U_{2,4},v)$. Let $(X_{i},i\in N_{4})$ be distributed on the rows of $\mathbf{T}\triangleq\left[\begin{matrix}\mathbf{T}_{0},\\ \dots\\ \mathbf{T}_{v-1}\end{matrix}\right]$ such that the probability mass of each row of $\mathbf{T}_{i}$ is $\frac{p_{i}}{v^{2}}$. Now we show the entropy function of $(X_{i},i\in N_{4})$ is $(a,b)$. By (III.4.24), $|\mathcal{X}_{4}|=v^{2}$ and each entry $j\in\mathcal{X}_{4}$ occurs only in one of the arrays $\mathbf{T}_{0}(4)$, $\dots$, $\mathbf{T}_{v-1}(4)$, which implies that for any subset $A$ such that $\{4\}\subseteq A\subseteq N_{4}$, each $x_{A}\in\mathcal{X}_{A}$ appears $\Delta_{A}$ times, where
$\Delta_{A}:=v^{\mathbf{r}_{U_{2,4}}(N_{4})-\mathbf{r}_{U_{2,4}}(A)},$
and $\mathbf{r}_{U_{2,4}}$ is the rank function of $U_{2,4}$. So
$H(X_{A})$ $=H\Big(\underbrace{\frac{p_{0}\Delta_{A}}{v^{2}},\dots,\frac{p_{0}\Delta_{A}}{v^{2}}}_{v^{\mathbf{r}_{U_{2,4}}(A)}},\dots,$
$\qquad\underbrace{\frac{p_{v-1}\Delta_{A}}{v^{2}},\dots,\frac{p_{v-1}\Delta_{A}}{v^{2}}}_{v^{\mathbf{r}_{U_{2,4}}(A)}}\Big)$ (III.4.25)
$=H(p_{0},p_{1},\dots,p_{v-1})+\log v\cdot\mathbf{r}_{U_{2,4}}(A)$ (III.4.26)
$=a+(a+b)\cdot\mathbf{r}_{U_{2,4}}(A)$ (III.4.27)
$=a\cdot(1+\mathbf{r}_{U_{2,4}}(A))+b\cdot\mathbf{r}_{U_{2,4}}(A).$ (III.4.28)
@@
$H(X_{3})$ $=H(\frac{v(p_{0}+p_{1}+\dots,+p_{v-1})}{v^{2}},\dots,$
$\frac{v(p_{0}.+p_{1}+\dots,+p_{v-1})}{v^{2}})$ (III.4.29)
$=\log{v}.$ (III.4.30)
@@
$H(X_{1},X_{3})$ $=H(X_{2},X_{3})=H(\frac{p_{0}+p_{1}+\dots,+p_{v-1}}{v^{2}},\dots,$
$\frac{p_{0}+p_{1}+\dots,+p_{v-1}}{v^{2}})$ (III.4.31)
$=2\log{v}.$ (III.4.32)
@@
$H(X_{1},X_{2},X_{3})$ $=H(\underbrace{\frac{p_{0}}{k^{2}},\frac{p_{0}}{v^{2}},\dots,\frac{p_{0}}{v^{2}}}_{v^{2}},\dots,$
$\qquad\underbrace{\frac{p_{v-1}}{v^{2}},\frac{p_{v-1}}{v^{2}},\dots,\frac{p_{v-1}}{v^{2}}}_{v^{2}})$ (III.4.33)
$=H(p_{0},\dots,p_{v-1})+2\log{v}$ (III.4.34)
$=a+2(a+b)=3a+2b.$ (III.4.35)
@@
$p_{1}=p_{4}=p_{6}=p_{7},$ (III.4.36)
$p_{2}=p_{3}=p_{5}=p_{8}.$ (III.4.37)
Assume that either (III.4.36) or (III.4.37) vanishes, then $\mathbf{T}_{0}$ degenerates a $\mathrm{VOA}(U_{2,3},2)$ $\mathbf{T}_{1}$ and $(X_{1},X_{2},X_{3})$ is uniformly distributed on the rows of $\mathbf{T}_{1}$. Together with (III.4.18) and (III.4.21),
$a=0,b=\log 2,$ (III.4.38)
which contradicts the fact that $\mathbf{h}=(0,\log 2)$ is non-entropic. Hence, both (III.4.36) and (III.4.37) must be positive. By (III.4.1), $X_{4}$ is a function of $(X_{1},X_{2},X_{3})$, which implies that $(X_{i},i\in N_{4})$ must be distributed on a $\mathbf{T}$ such that $\mathbf{T}(1,2,3)=\mathbf{T}_{0}$. By (III.4.15) and (III.4.16), for each $x_{4}\in\mathcal{X}_{4},x_{j}\in\mathcal{X}_{j}$, $(x_{j},x_{4})$ appear exactly once on the row of $\mathbf{T}(j,4)$ for $j=1,2,3$. Additionally, $X_{4}$ is independent of $X_{i},i=1,2,3$ by (III.4.7), (III.4.9), and (III.4.10). There exists a unique $\mathbf{T}$ satisfying the above information equalities up to symmetry, and
@@
Now we show an inner bound on the entropy region on $F$ that $a+b=\log 6$ and $\log 2\leq a\leq\log 6$. Let $p_{i}>0$, $i=0,1,2$, and $p_{0}+p_{1}+p_{2}=1$ and $H(p_{0},p_{1},p_{2})=a-\log 2$. Let $\mathbf{T}^{(k)}_{0},k=1,2$ be arrays defined as follows.
@@
$\mathbf{T}^{(k)}_{i}(1,2)$ $=\mathbf{T}^{(k)}_{1}(1,2),$ (III.4.44)
$\mathbf{T}^{(k)}_{i}(3;j)$ $=\sigma_{i}(\mathbf{T}^{(k)}_{1}(3;j))\text{ for }j\in N_{36}$ (III.4.45)
$\mathbf{T}^{(k)}_{i}(4;j)$ $=\mathbf{T}^{(k)}_{1}(4;j)+12i\text{ for }j\in N_{36},$ (III.4.46)
@@
$H(X_{4})$ $=H(\underbrace{\frac{p_{0}}{12},\frac{p_{0}}{12},\dots,\frac{p_{0}}{12}}_{12},\dots,$ (III.4.47)
$\qquad\underbrace{\frac{p_{2}}{12},\frac{p_{2}}{12},\dots,\frac{p_{2}}{12}}_{12})$ (III.4.48)
$=H(p_{0},p_{1},p_{2})+\log{12}$ (III.4.49)
$=a-\log 2+\log 12$ (III.4.50)
$=2a+b.$ (III.4.51)
@@
$p_{X_{1}}(j)=\frac{12p_{0}}{72}+\frac{12p_{1}}{72}+\frac{12p_{2}}{72}=\frac{1}{6},$ (III.4.52)
@@
$H(X_{1},X_{4})$ $=H(X_{1},X_{2},X_{3},X_{4})$ (III.4.55)
$=H(\underbrace{\frac{p_{0}}{72},\frac{p_{0}}{72},\dots,\frac{p_{0}}{72}}_{72},\dots,$ (III.4.56)
$\qquad\underbrace{\frac{p_{2}}{72},\frac{p_{2}}{72},\dots,\frac{p_{2}}{72}}_{72})$ (III.4.57)
$=H(p_{0},p_{1},p_{2})+\log{72}$ (III.4.58)
$=a-\log 2+\log 72$ (III.4.59)
$=3a+2b.$ (III.4.60)
@@
$p_{X_{1}X_{3}}(x_{1},x_{2})=\frac{2p_{0}}{72}+\frac{2p_{1}}{72}+\frac{2p_{2}}{72}=\frac{1}{36},$ (III.4.65)
@@ III-E Discussion
In this section, we characterized $10$ types of $2$-dimensional faces of $\Gamma_{4}$. For a face $F$ with one of its two extremes rays containing a rank-1 matroid, we can see that entropy region $F^{*}$ has a similar shape to the entropy region on $(U_{2,3},U^{12,3}_{1,2})$ characterized by Matúš in [10] or $(U_{2,3},U^{1,3}_{1,1})$ by Chen and Yeung in [11]. The characterization of such faces are generalized to faces of general $n$ and both extreme rays containing matroids in [32].
As each $\mathrm{MVOA}(P,v)$ with $M$ a rank 2 or greater integer polymatroid corresponds to a type of orthogonal Latin hypercubes, the characterization of a face with both extreme rays containing an integer polymatroid exceeding 1, in Subsection III-B-III-D, breeds a new combinatorial structure which can be considered as an intermediate form of the two types of orthogonal Latin hypercubes. Specifically, in Subsection III-C, three faces are characterized by the Latin square decompositions, which can be considered as three new types of Latin square substructures.